In a chemistry class, students were learning about boiling point of various elements.Reuben has compiled a table that gives boiling points of each of various elements.
The lowest boiling point is of Element X: – 200 degrees Fahrenheit.
The highest boiling point is of Element Y: 550 degrees Fahrenheit
Find the difference between the lowest and highest boiling point.
I know how to calculate the difference.
-200 – 550 = -750.
I am not sure if the answer is written as a negative or positive. Isn’t difference in temperature always written as its absolute value?
Will the answer be – 750 F or 750 F ?
Would greatly appreciate if you could explain what the correct answer is.

Found 2 solutions by math_helper, MathTherapy:
Answer by math_helper(2461)


The way I interpret the question, you can answer it either way. However, when answering, you should take the sign into account and specify a magnitude difference and whether it is LOWER or HIGHER:

-200F - 550F = -750F —> thus "Element X boils at a temperature that is 750F lower than that of element Y"

550F - (-200F) = 750F —> thus "Element Y boils at a temperature that is 750F higher than that of element X"

Both of these answers say the same thing, just with different points of reference.

Answer by MathTherapy(10555)


In a chemistry class, students were learning about boiling point of various elements.Reuben has compiled a table that gives boiling points of each of various elements.
The lowest boiling point is of Element X: – 200 degrees Fahrenheit.
The highest boiling point is of Element Y: 550 degrees Fahrenheit
Find the difference between the lowest and highest boiling point.
I know how to calculate the difference.
-200 – 550 = -750.
I am not sure if the answer is written as a negative or positive. Isn’t difference in temperature always written as its absolute value?
Will the answer be – 750 F or 750 F ?
Would greatly appreciate if you could explain what the correct answer is
To find the difference between the HIGHEST and the LOWEST boiling points, you just SIMPLY SUBTRACT the LOWER of the 2 temperatures from the HIGHER of the 2 temperatures. 
This gives us: